# Sparsing Law: Towards Large Language Models with Greater Activation Sparsity

### v1

@@ 3.3 Setting of Scaling Analysis
We mainly focus on the activation sparsity of foundation models, which only undergo the pre-training stage. However, before we evaluate models on task-specific benchmarks, we follow MiniCPM (Hu et al., 2024) to conduct a decay stage, where instruction-tuning data is additionally added for training. Thereby, we can obtain more reasonable results on benchmarks. Besides, considering the influence of various training hyper-parameters, we follow the optimal batch sizes, optimal learning rates, and the WSD learning rate scheduler of MiniCPM (Hu et al., 2024).
Evaluation Settings
@@
We introduce a tiny validation dataset and two groups of benchmarks, including commonsense reasoning and reading comprehension, for evaluation. We also test our model on more complex coding or knowledge benchmarks but fail to obtain a performance significantly surpassing the random level (See Appendix H), mainly due to the small parameter scales. For the measurement of sparsity, to eliminate the impact of stochastic factors (especially the sparsity fluctuations during the early stage), we use a sparsity stabilizing strategy to obtain smoother sparsity-data curves (see Appendix B for more details).
@@
If there are no special statements, the training loss, validation loss, and perplexity are all calculated on models that only complete pre-training (the latter two metrics are computed on the validation data), and the task-specific performance is evaluated on checkpoints after the decay stage. See Appendix E and F for more details about datasets and training settings, respectively.
4 Experiments
4.1 Rationality of PPL-$p\%$ Sparsity
As stated in Section 1, a rational metric for activation sparsity should have: versatility across model architectures, performance-awareness, and the precise recognition of weakly-contributed neurons. While the former two properties are already demonstrated in Section 1, we mainly discuss whether PPL-$p\%$ sparsity has the third one in this section. Specifically, PPL-$p\%$ sparsity should wisely recognize more weakly-contributed neurons while minimizing performance degradation simultaneously. In other words, our metric should strike a better trade-off between performance and sparsity. We introduce the following baselines for comparison:
(1) Straightforward ReLU: The most simple setting that uses the zero threshold and is only applicable to ReLU: $\mathcal{D}=\{i|s_{i}=0\}$.
(2) Top-$k$ sparsity, widely adopted in the MoE architectures (Fedus et al., 2022; He, 2024), enforces each layer to consistently maintain $k$ activated neurons, whose absolute values of activation scores rank in the top-k ones among all the neurons of that layer. Obviously, we have $|\mathcal{D}|=d_{f}-k$, and the Top-$k$ sparsity method holds a constant sparsity ratio across all layers.
(3) FAT-$\epsilon$ sparsity (Kurtz et al., 2020) (FAT denotes forced activation threshold) similarly introduces a global hyper-parameter $\epsilon$ as the threshold shared by all layers, namely $\mathcal{D}=\{i||s_{i}|<\epsilon\}$. Note that this is slightly different from the original FATReLU by Kurtz et al. (2020) as we compute the absolute values of activation scores to accommodate SiLU.
Figure 3: The PPL-activation Pareto curve of our PPL-$p\%$ sparsity versus two baselines within models of different scales. “Straightforward ReLU” is only applicable to ReLU-activated models.
@@ 4.1 Rationality of PPL-$p\%$ Sparsity
Table 1: The average evaluation scores (%) on two task groups, where C.R. refers to commonsense reasoning and R.C. refers to reading comprehension. The second column represents settings with different $p\%$ values, with “dense” indicating the most accurate case where $p=0$.
0.1B 0.2B 0.4B 0.8B 1.2B
@@
C.R. dense 49.6 49.5 52.0 52.2 54.7 55.8 56.8 57.6 60.0 59.6
PPL-1% 49.1 49.9 51.7 52.4 54.6 55.8 55.9 57.6 59.6 59.6
PPL-5% 49.2 49.0 51.7 52.0 54.3 55.1 56.3 57.1 59.3 58.8
PPL-10% 49.4 48.7 51.6 51.9 54.9 55.2 55.6 56.4 59.3 59.3
R.C. dense 28.2 27.7 40.7 40.2 44.0 41.8 44.8 43.3 53.2 54.8
PPL-1% 28.4 28.0 39.7 39.6 42.9 40.9 43.2 44.3 53.3 55.4
PPL-5% 26.9 26.5 38.6 36.8 40.8 38.2 42.2 40.7 53.3 52.6
PPL-10% 26.2 24.8 38.6 34.4 39.9 35.3 40.3 38.8 52.9 51.1
As demonstrated by Figure 3, PPL-$p\%$ sparsity obtains the best trade-off between sparsity and performance, with the lowest PPL given a target sparsity ratio. This is brought by its wise recognition of inactivated neurons through adaptive thresholds. Moreover, the performance-aware definition of PPL-$p\%$ sparsity makes us explicitly grasp the sparsity ratio at each desired perplexity.
Another problem lies in how the hyper-parameter $p\%$ influences task-specific performance. To inspect this, we evaluate models of different scales on the benchmarks in Section 3.3 and tune their sparsity through distinct $p\%$ values. As shown in Table 1, with the value of $p\%$ increasing (intrinsically promoting greater sparsity), the reading comprehension performance is considerably impaired, corresponding to the trade-off between sparsity and performance. Notably, in both task groups, the average performance of “PPL-1%” is comparable to that of the theoretically most accurate “dense” setting. Therefore, we assume PPL-1% sparsity as a reliable performance-unimpaired metric and employ it to compute sparsity in the following discussions.
4.2 Activation Sparsity with the Amount of Training Data and Activation Functions
To obtain the scaling relationship between the activation sparsity and the amount of training data, we pre-train models with different numbers of parameters and two activation functions (i.e., ReLU and SiLU), respectively, and then evaluate the sparsity level of their checkpoints using PPL-$1\%$. After careful attempts, we find that the curve of activation ratios to the amount of training data is easier to fit than that of the sparsity ratio. Therefore, we will frequently use activation ratios instead of sparsity ratios in the following sections, whose trend is plotted in Figure 4.
@@ 4.2 Activation Sparsity with the Amount of Training Data and Activation Functions
Figure 4: The trend of activation ratios (hereinafter using PPL-$1\%$ sparsity) of models with different scales and activation functions during the pre-training stage. The fitted curves are plotted in brown. The number of training tokens is no less than 190 times the scale of non-embedding parameters.
@@
For ReLU models, we observe a logspace power-law relationship between the activation ratio $A_{ReLU}(D)$ and the amount of training data $D$, expressed in the following formula:
@@
where $A_{0}>0$ is the limit activation ratio with infinite training data and we have $c,\alpha>0$. Obviously, this is a convergent decreasing function, indicating that more training data can potentially make ReLU models more sparsely activated.
@@
where similarly, $A_{0}>0$ is the limit activation ratio and $c,\alpha>0$. Note that this is a convergent increasing function, and thus more training data will impair the activation sparsity of SiLU models. See Appendix D for the algorithm of curve fitting and the results (i.e., coefficients).
@@
As for the selection of activation functions, by comparing the sparsity dynamics, we can conclude that the activation sparsity achieved by ReLU is significantly greater than that of SiLU. Besides, the task-specific performance in Table 1 and the trend of training loss in Appendix A reveal the comparable performance between ReLU and SiLU activations. Based on the above observations, ReLU is more competent as the activation function than SiLU due to three advantages: an increasing trend of sparsity, significantly higher sparsity ratio, and comparable performance.
4.3 Activation Sparsity with the Width-Depth Ratio
@@
The width-depth ratio, defined as the ratio of the hidden dimension to the layer number, reflects the shape of a Transformer and is a key architectural property that potentially influences activation sparsity. To inspect its influence on the activation sparsity, we conduct experiments on the 0.1B ReLU-activated model and select 9 different width-depth ratios. The limit activation ratio and the limit training loss are plotted in Figure 6 and Figure 6 respectively.
@@
As demonstrated by Figure 6, under a bottleneck point (about 114 for 0.1B), the activation ratio linearly increases with the width-depth ratio. However, given a width-depth ratio greater than this bottleneck, the activation ratio fluctuates around 8%. From the sparsity aspect, a smaller width-depth ratio is definitely more helpful. However, the dynamics shown in Figure 6 demonstrate that there exists the best interval of width-depth ratio for the lowest training loss (from 74 to 282 for 0.1B). Therefore, to maintain the best performance while promoting greater activation sparsity, the best width-depth ratio should fall on the smallest point of this interval (i.e., around 74 for 0.1B).
Figure 5: The limit activation ratios on 0.1B ReLU-activated models.
Figure 6: The limit training loss on 0.1B ReLU-activated models.
Figure 7: The limit activation ratio for pre-trained models with different scales and activation functions.
Figure 8: The derivative trends of the sparsity-data curve with the increase of data-scale ratio, within ReLU/SiLU models of distinct scales.
4.4 Activation Sparsity with the Parameter Scale
@@
To obtain comprehensive scaling properties of activation sparsity with the increase of scales (i.e., the number of non-embedding parameters), we obtain the limit activation ratio of the above pre-trained models with 5 distinct scales but similar width-depth ratios. From the results plotted in Figure 8, we can reach the first observation that under similar width-depth ratios, the limit activation ratio as the amount of training data approaches infinity is weakly related to the parameter scale. For SiLU settings, the activation ratio decreases slightly by 2.7 points from 0.1B to 1.2B. By contrast, for ReLU settings, the activation ratio marginally increases by 1.7 points from 0.1B to 1.2B.
@@
To reflect the evolving dynamics of sparsity, we compute the derivatives of the sparsity-data curve as fitted in Section 4.2 and plot the trend of derivatives with the increase of data-scale ratio [^3] The data-scale ratio means the ratio of the number of training tokens to the parameter scale. We choose this variable as previous works have demonstrated the roughly proportional relationship between the optimal amount of training data and the parameter scale (Hoffmann et al., 2022; Besiroglu et al., 2024).. The results in Figure 8 clearly demonstrate that smaller models tend to converge faster than larger models to the limit, as the absolute values of their derivatives are much larger. We try to explain the above observations as follows.
@@
Figure 9: The distribution of the neuron activation frequencies within models of distinct scales. Four datasets from the pre-training data are involved.
Figure 10: The activation ratio (%) distributions of 71,549 tokens sampled from the vocabulary. We conduct a pair-wise comparison of the average activation ratio of each token within models of different scales. Note that the red line is the $y=x$ curve.
To inspect the dataset-wise activation distributions under different parameter scales, we consider four datasets, each a subset of the pre-training data: Code, Wikipedia, Math, and Chinese. Next, we compute the distribution of activation frequencies (i.e., the times that a neuron is activated divided by the total number of tokens) among the neurons within models of different scales. As demonstrated by Figure 9, for all the datasets, the distribution patterns of neuron activation frequencies are similar across different scales. While this observation holds on average, special cases exist in certain layers (see Appendix G).
@@
From different aspects, the above two experiments both support the insensitiveness of the activation pattern to the parameter scale. This can potentially provide one explanation for why the activation sparsity is quite weakly correlated with the model sizes.
@@
Deduction: Smaller models converge faster to the limit activation ratio mainly due to their small amount of neurons. To simplify this problem, we model the specialization of neurons as a grouping process, where each neuron can be placed into zero or more groups (considering the potential existence of dead neurons and versatile neurons). Suppose the $d_{f}$ neurons should specialize into $G$ groups, each of them having $t_{1},t_{2},...,t_{G}$ neurons respectively. Based on the assumption of similar activation patterns and neuron specialization, the ratio of neurons placed in each group (i.e., $0<t_{i}/d_{f}\leq 1,\ i=1,2,...,G$) should be shared across different parameter scales. We can obtain the number of all the possible grouping results $T(d_{f})$ easily,
@@
where $C_{d_{f}}^{t_{i}}$ is the combinatorial number, the number of possibilities to select $t_{i}$ neurons from $d_{f}$ ones. Obviously, $T(d_{f})$ grows in a factorial speed with $d_{f}$, much faster than the linear function. For larger models, the number of neuron specialization possibilities is significantly greater than that of smaller ones. Therefore, it takes more training expenses for larger models to form stable neuron specialization and approach the limit activation ratio.

### v2

@@ 3 Quantitative Study Settings
Training settings We mainly explore the activation sparsity of foundation models, which only undergo the pre-training stage. However, before we evaluate models on task-specific benchmarks, we follow recent LLMs (Dubey et al., 2024; Hu et al., 2024) to conduct a decay stage, where instruction-tuning data is added. Thereby, we can obtain more reasonable results on benchmarks. We also follow the optimal batch sizes and learning rates in recent LLMs.
@@
Evaluation settings We introduce a tiny validation dataset and two groups of benchmarks for evaluation, including commonsense reasoning (C.R.) and reading comprehension (R.C.). For the measurement of sparsity, to eliminate the impact of stochastic factors (especially the sparsity fluctuations during the early stage), we employ a sparsity stabilizing strategy (see Appendix E).
@@
If there are no special statements, the training loss, validation loss, and perplexity (on validation data) are all calculated on models that only complete pre-training, and the task-specific performance is evaluated on checkpoints after the decay stage. See Appendix HI, and K for more details about datasets and training settings.
4 How can activation sparsity be measured more accurately?
4.1 Recognition of Weakly-Contributed Neurons
@@
Table 1: The average evaluation scores (%) on two task groups. The second column represents settings with different PPL increase tolerances $p\%$, with “Dense” indicating the most accurate case where $p=0$. The marker “$\Delta$” means the score difference to the corresponding dense setting.
0.1B 0.2B 0.4B 0.8B 1.2B Avg.
@@
C.R. Dense 49.6 49.5 52.0 52.2 54.7 55.8 56.8 57.6 60.0 59.6 54.78
$\Delta$ CETT-PPL-1% --0.5 $+$0.4 --0.3 $+$0.2 --0.1 0 --0.9 0 --0.4 0 --0.16
$\Delta$ CETT-PPL-5% --0.4 --0.5 --0.3 --0.2 --0.4 --0.7 --0.5 --0.5 --0.7 --0.8 --0.50
$\Delta$ CETT-PPL-10% --0.2 --0.8 --0.4 --0.3 $+$0.2 --0.6 --0.2 --1.2 --0.7 --0.3 --0.45
R.C. Dense 28.2 27.7 40.7 40.2 44.0 41.8 44.8 43.3 53.2 54.8 41.87
$\Delta$ CETT-PPL-1% $+$0.2 $+$0.3 --1.0 --0.6 --1.1 --0.9 --1.6 $+$1.0 $+$0.1 $+$0.6 --0.30
$\Delta$ CETT-PPL-5% --1.3 --1.2 --2.1 --3.4 --3.2 --3.6 --2.6 --2.6 $+$0.1 --2.2 --2.21
$\Delta$ CETT-PPL-10% --2.0 --2.9 --2.1 --5.8 --4.1 --6.5 --4.5 --4.5 --0.3 --3.7 --3.64
@@
Figure 4: The trend of activation ratios of models with different parameter numbers and activation functions during the pre-training stage. The fitted curves are plotted in brown.
As stated above, the crucial part of a metric for activation sparsity is the accurate recognition of weakly-contributed neurons. In this section, we compare the following four methods to determine weakly-contributed neurons in FFNs (Refer to Eq. (2) for the definition of neurons):
(1) Straightforward ReLU is the most simple but commonly used setting, which uses a zero threshold and is only applicable to ReLU: $\mathcal{D}=\{i|s_{i}=0\}$.
(2) Top-$k$, widely adopted in the MoE architectures (Fedus et al., 2022), enforces each layer to consistently maintain $k$ activated neurons, whose absolute activation values rank in the top-k ones among all the neurons of that layer. Obviously, we have $|\mathcal{D}|=d_{f}-k$, and the Top-$k$ sparsity method holds a constant sparsity ratio across all layers.
(3) FAT-$\epsilon$ (Kurtz et al., 2020) (FAT denotes forced activation threshold) similarly introduces a global hyper-parameter $\epsilon$ as the threshold shared by all layers, namely $\mathcal{D}=\{i||s_{i}|<\epsilon\}$. Note that this is slightly different from the original FATReLU by Kurtz et al. (2020) as we compute the absolute values of activation values to accommodate SiLU.
(4) CETT (Zhang et al., 2024a), as introduced in Section 2.2, requires each layer to share the same $L_{2}$ norm relative error after weakly-contributed neurons are skipped in calculation, while the layer-wise activation thresholds are adaptively searched accordingly.
As demonstrated by Figure 2 and Figure 3, CETT obtains the best trade-off between sparsity and performance. Under the same target sparsity ratio, CETT can always achieve the lowest PPL, and a comparable or higher evaluation score on downstream tasks. These indicate the highest accuracy of CETT in recognizing weakly-contributed neurons.
4.2 Introduction of PPL Increase Tolerance
Despite the advantages of CETT, it remains a problem how to choose the best hyper-parameter, namely the shared $L_{2}$ norm relative error. Therefore, we introduce CETT-PPL-$p\%$, which denotes the sparsity ratio measured by CETT when the PPL on validation data rises by $p\%$ compared with the dense setting (i.e., with all neurons activated). Given a PPL increase tolerance of $p\%$, we can conduct a binary search to find the hyper-parameter that just meets the tolerance. The search algorithm is described in Appendix F.
To find a proper $p\%$ tolerance, we inspect its impact on downstream task performance. As shown in Table 1, with PPL increasing (intrinsically promoting greater sparsity), the reading comprehension performance is considerably impaired, corresponding to the trade-off between sparsity and performance. Notably, in both task groups, the average performance of “CETT-PPL-1%” is comparable to that of the theoretically most accurate “Dense” setting. Therefore, we assume CETT-PPL-1% sparsity as a reliable performance-friendly metric and employ it to compute sparsity in the following discussions.
Figure 5: Limits of activation ratios on 0.1B ReLU models.
Figure 6: Limit training loss on 0.1B ReLU models.
Figure 7: The limits of activation ratios on models with different scales and activation functions.
Figure 8: The derivative trends of the sparsity-data curve with the increase of data-scale ratio.
5 How is activation sparsity affected by the model architecture and training process?
5.1 Amount of Pre-Training Data and Choice of Activation Functions
To obtain the scaling relationship between the activation sparsity and the amount of pre-training data, we pre-train models with different parameter numbers and two activation functions (i.e., ReLU and SiLU), respectively, and then evaluate the sparsity level of their checkpoints using CETT-PPL-$1\%$. After careful attempts, we find that the curve of activation ratios to the amount of pre-training data is easier to fit than that of sparsity ratios. Therefore, we will frequently use activation ratios instead of sparsity ratios in the following sections, whose trend is plotted in Figure 4.
@@ 5.1 Amount of Pre-Training Data and Choice of Activation Functions
For ReLU models, we observe a logspace power-law relationship between the activation ratio $A_{ReLU}(D)$ and the amount of pre-training data $D$, expressed as follows:
@@
where $A_{0}>0$ is the limit of activation ratio with infinite training data and $c,\alpha>0$. This is a convergent decreasing function, indicating that more training data can potentially make ReLU models more sparsely-activated.
@@
where similarly, $A_{0}>0$ is the limit of activation ratio and $c,\alpha>0$. Note that this is a convergent increasing function, and thus more training data will impair the activation sparsity of SiLU models. See Appendix G for the algorithm of curve fitting and the results (i.e., coefficients).
@@
As for the selection of activation functions, by comparing the sparsity dynamics, we can conclude that the activation sparsity achieved by ReLU is significantly greater than that of SiLU. Besides, the task-specific performance in Table 1 and the trend of training loss in Appendix D reveal the comparable performance between ReLU and SiLU activations. Based on the above observations, ReLU is more competent as the activation function than SiLU due to three advantages: an increasing trend of sparsity, significantly higher sparsity ratio, and comparable performance.
5.2 Width-Depth Ratio
@@
The width-depth ratio, defined as the ratio of the hidden dimension to the layer number, reflects the shape of a Transformer and is a key architectural property that potentially influences activation sparsity. To inspect its influence on the activation sparsity, we conduct experiments on the 0.1B ReLU-activated model and select 9 different width-depth ratios. The limits of the activation ratio and the training loss are plotted in Figure 6 and Figure 6 respectively.
@@
As demonstrated by Figure 6, under a bottleneck point (about 114 for 0.1B), the activation ratio linearly increases with the width-depth ratio. After this bottleneck, the activation ratio fluctuates around 8%. From the sparsity aspect, a smaller width-depth ratio is more helpful. However, Figure 6 demonstrates that an extremely small width-depth ratio causes significant performance degradation. This is attributed to the training instability of deep Transformers (Petty et al., 2023; Wu & Tang, 2024). Therefore, the best width-depth ratio should fall on the smallest point of the interval that ensures stable training and the best performance (e.g., from 74 to 282 for 0.1B).
5.3 Parameter Scale
Figure 9: The distribution of the neuron activation frequencies on four datasets within models of distinct scales.
Figure 10: The activation ratio (%) distributions of 71,549 tokens sampled from the vocabulary. We conduct a pair-wise comparison of the average activation ratio of each token within models of different scales. The red line is the $y=x$ curve.
@@ 5.3 Parameter Scale
To obtain comprehensive scaling properties of activation sparsity with the increase of scales (i.e., the number of non-embedding parameters), we obtain the limit of activation ratio on the above pre-trained models with 5 distinct scales but similar width-depth ratios. From the results plotted in Figure 8, we can reach the first observation that under similar width-depth ratios, the limit of activation ratio, as the amount of pre-training data approaches infinity, is weakly related to the parameter scale. For SiLU settings, the activation ratio decreases slightly by 2.7 points from 0.1B to 1.2B. By contrast, for ReLU settings, the activation ratio marginally increases by 1.7 points from 0.1B to 1.2B.
@@
To reflect the evolving dynamics of sparsity, we compute the derivatives of the sparsity-data curve as fitted in Section 5.1 and plot the trend of derivatives with the increase of data-scale ratio [^2] The data-scale ratio means the ratio of the number of training tokens to the parameter scale. We choose this variable as previous works have demonstrated the roughly proportional relationship between the optimal amount of pre-training data and the parameter scale (Hoffmann et al., 2022; Besiroglu et al., 2024).. The results in Figure 8 clearly demonstrate that smaller models tend to converge faster than larger models to the limit, as the absolute values of their derivatives are much larger. We try to explain the above observations.
@@
To inspect the dataset-wise activation distribution, we consider four subsets of the pre-training data and compute the distribution of activation frequencies (i.e., the times that a neuron is activated divided by the total number of tokens) among the neurons within models of different scales. As demonstrated by Figure 9, for all datasets, the distribution patterns of neuron activation frequencies are similar across different scales. While this observation holds on average, special cases exist in certain layers (see Appendix J).
@@
The above two experiments both support the insensitivity of the activation pattern to the parameter scale. This can potentially provide one explanation for why the activation sparsity is quite weakly correlated with the model sizes.
@@
Deduction: Smaller models converge faster to the limit of activation ratio mainly due to their small amount of neurons. To simplify this problem, we model the specialization of neurons as a grouping process, where each neuron can be placed into zero or more groups (considering the potential existence of dead neurons and versatile neurons). Suppose the $d_{f}$ neurons should specialize into $G$ groups, each of them having $t_{1},t_{2},...,t_{G}$ neurons respectively. Based on the assumption of similar activation patterns and neuron specialization, the ratio of neurons placed in each group (i.e., $0<t_{i}/d_{f}\leq 1,\ i=1,2,...,G$) should be shared across different parameter scales. We can obtain the number of all the possible grouping results $T(d_{f})$ easily,
@@
where $C_{d_{f}}^{t_{i}}$ is the combinatorial number, the number of possibilities to select $t_{i}$ neurons from $d_{f}$ ones. Obviously, $T(d_{f})$ grows in a factorial speed with $d_{f}$, much faster than the linear function. For larger models, the number of neuron specialization possibilities is significantly greater, and thus more training expenses are required to form stable neuron specialization and approach the limit of activation ratio.
6 How can we build a more sparsely activated and efficient LLM?
Figure 11: The activation-data trend of the 2.4B ReLU model, with the fitted logspace power-law presented.
Based on the above findings, we finally come to the approach to building an LLM with greater activation sparsity: Use ReLU as the activation function with a larger amount of pre-training data, and a small width-depth ratio within the interval ensuring the training stability.
To validate the reliability and generalizability of our findings, we train a 2.4B model based on the above experience. Equipped with ReLU activation, near 800B training data, and a width-depth ratio close to the above models (also small enough and close to the smallest point of the training stability interval), this model achieves a low limit activation ratio of 6.48%, quite close to those values of models from 0.1B to 1.2B. Moreover, its activation-data trend can also be well fitted with a decreasing logspace power-law, as shown in Figure 11. These are all consistent with previous findings.
Moreover, to demonstrate the practical acceleration value, we compare the decoding speed of the 2.4B model with PowerInfer (Song et al., 2023) and “llama.cpp” (Gerganov, 2023), respectively. While PowerInfer utilizes sparsity to save computation, “llama.cpp” only conducts dense FFN computation. Consequently, the former achieves 4.1$\times$ speedup compared with the latter, revealing the potential of activation sparsity for efficient LLMs. Refer to Appendix L for experiment details.